# When to Trust LLMs: Aligning Confidence with Response Quality and Exploring Applications in RAG

### v1

@@ Abstract
Despite the success of large language models (LLMs) in natural language generation, much evidence shows that LLMs may produce incorrect or nonsensical text. This limitation highlights the importance of discerning when to trust LLMs, especially in safety-critical domains. Existing methods, which rely on verbalizing confidence to tell the reliability by inducing top-k responses and sampling-aggregating multiple responses, often fail, due to the lack of objective guidance of confidence. To address this, we propose CONfidence-Quality-ORDer-preserving alignment approach (CONQORD), leveraging reinforcement learning with a tailored dual-component reward function. This function encompasses quality reward and order-preserving alignment reward functions. Specifically, the order-preserving reward incentivizes the model to verbalize greater confidence for responses of higher quality to align the order of confidence and quality. Experiments demonstrate that our CONQORD significantly improves the alignment performance between confidence levels and response accuracy, without causing the model to become over-cautious. Furthermore, the aligned confidence provided by CONQORD informs when to trust LLMs, and acts as a determinant for initiating the retrieval process of external knowledge. Aligning confidence with response quality ensures more transparent and reliable responses, providing better trustworthiness.
When to Trust LLMs: Aligning Confidence with Response Quality and Exploring Applications in RAG
Shuchang Tao♠,♡ Liuyi Yaothanks: Corresponding author  Hanxing Ding Yuexiang Xie  Qi Cao Fei SunJinyang GaoHuawei ShenBolin Ding Alibaba Group {taoshuchang.tsc, yly287738, yuexiang.xyx, jinyang.gjy, bolin.ding}@alibaba-inc.com CAS Key Laboratory of AI Safety, Institute of Computing Technology, Chinese Academy of Sciences {dinghanxing18s, caoqi, sunfei, shenhuawei}@ict.ac.cn
@@ 1 Introduction
Large Language Models (LLMs) have excelled in natural language understanding and generation, finding use in areas such as healthcare, fraud detection, and legal assistance  Brown et al. (2020); Anil et al. (2023); American Psychological Association (1983). However, mounting evidence suggests that LLMs can generate incorrect or non-sense text, such as made-up citations or inaccurate medical information, risking errors in critical applications  Ji et al. (2023); Zhang et al. (2023b); Agrawal et al. (2023); Du et al. (2023); Cohen et al. (2023). The urgent question is: When can we trust LLMs? Addressing this issue is crucial for preventing users from uncritically accepting misleading information. Furthermore, resolving this problem helps determine when to trust LLMs and when to seek external knowledge.
@@
Figure 1: Absence or misalignment of confidence and response. Vanilla LLMs may generate bad responses and are unable to generate confidence. Existing methods include verbalizing confidence levels (highlighted in orange) in the output to indicate the model’s uncertainty, yet they may still provide bad responses with overly high confidence, revealing a misalignment between expressed confidence and actual response quality.
Confidence calibration, which ensures that the probability associated with a predicted class label accurately reflects its likelihood of being correct, is critical for the trustworthiness of machine learning systems Guo et al. (2017). Proper calibration of these expressions is imperative for the models’ reliability, transparency, and interpretability. In traditional classification tasks, the predicted probabilities should reflect the actual likelihood of correctness based on ground truth label Guo et al. (2017); Minderer et al. (2021). Conversely, in generative tasks, calibrating confidence is notably challenging due to the absence of clear ground truth, presenting significant hurdles.
Recently, researchers have proposed methods to request LLMs to express confidence within the output text, either as numerical values or through qualitative descriptors of uncertainty Tian et al. (2023). Existing methods either retrieve the top-k responses to calibration or aggregate the confidence of sampling multiple responses Xiong et al. (2023). They rely on some heuristic assumptions, such as multi-samples should be consistent, and recalling top-k samples helps to calibrate confidence. However, as shown in Figure 1, these methods frequently lead to a discrepancy between the expressed confidence and the actual quality of the responses provided. Therefore, these methods may still yield incorrect responses accompanied by inappropriately high confidence, indicating a misalignment between the model’s confidence and its quality.
The core issue lies in the absence of a gold standard for confidence that aligns with response quality, thereby guiding the model to optimize effectively.
In this paper, we consider a Reinforcement Learning (RL) framework to address this issue, leveraging RL’s flexibility brought by various rewards functions Guo et al. (2021); Ziegler et al. (2019).
A straightforward reward strategy is to assign high scores to well-aligned response-confidence pairs while assigning low scores to misaligned ones, however, which can inadvertently encourage language models to take shortcuts, (empirically demonstrated in Section 3). Such a strategy might result in a preference for generating responses of inferior quality accompanied by low confidence levels. To tackle this, careful consideration must be given to the design of the reward function to ensure it promotes the generation of high-quality, well-aligned responses.
Based on the aforementioned insights, we propose a CONfidence-Quality-ORDer-preserving alignment approach, called CONQORD (sounds as Concord). We decouple the above reward into two components: (i) For response quality, we train a quality reward model that rates the response. (ii) For confidence alignment, we design an order-preserving criterion that maintains a consistent ordinal relationship between the verbalized confidence and response quality rating. We propose an alignment reward function rewarding consistency and penalizing discrepancies in rankings between samples. It encourages careful self-calibration and adapts dynamically to different contexts, while this order-preserving reward is less affected by outliers.
With this dual-component reward function, comprising both the quality reward model and the alignment reward function, we employ Proximal Policy Optimization (PPO) algorithm to align the verbalized confidence with the quality of the response, without causing the model to become over-cautious.
@@
Our experimental results demonstrate that this technique substantially improves the congruence between confidence alignment levels and the accuracy of responses without inducing excessive caution. Moreover, we evaluate the practicality of CONQORD’s calibrated confidence in downstream tasks, particularly within Retrieval-Augmented Generation (RAG) frameworks Ding et al. (2024); Wei et al. (2024a, b). Here, confidence scores inform the initiation of retrieval operations. Our experiments confirm that CONQORD’s confidence alignment reliably dictates the trustworthiness of LLM outputs, guiding the activation of external knowledge retrieval. The aligned confidence provided by CONQORD makes the responses more transparent and reliable.
@@ 2 Related Works
Confidence, or uncertainty, refers to the degree of certainty or assurance that accompanies a prediction or decision made by a model Geng et al. (2023). The calibration of confidence is essential for the reliability of machine learning systems, as it ensures that predicted probabilities match the true likelihood of outcomes as closely as possible Guo et al. (2017); Minderer et al. (2021). While in traditional classification tasks, this involves aligning predicted probabilities with actual ground truth labels, the task becomes more challenging for generative models due to the inherently ambiguous nature of the ground truth Gawlikowski et al. (2021); Liu et al. (2023); Ding et al. (2023).
@@ 3 Confidence Alignment via RL
We adopt a Reinforcement Learning (RL) framework to tackle the challenge of lacking a ground-truth standard for confidence assessment. Different from Supervised Fine-Tuning (SFT), which is dependent on labeled data, RL offers a more adaptable solution by allowing any indicator as a reward. We follow previous studies Ramamurthy et al. (2023); Wu et al. (2023), regard text language generation as a Markov Decision Process (MDP) while the remaining elements of MDP are listed in Appendix A. In the rest of this section, we elaborate on the reward strategy.
@@ 3.1 Preliminary Alignment Approach
To align the quality of responses with corresponding confidence levels, we propose a simplistic reward strategy as a preliminary approach utilizing Reinforcement Learning from Human Feedback (RLHF) with a reward model that assesses the degree of alignment. We refer to this preliminary approach as PreApproach.
@@ 3.2 Results and Discussions
The experimental results are illustrated in Figure 2 (a), which shows that PreApproach attains a lower Expected Calibration Error (ECE) Guo et al. (2017) than LLAMA-2 7b, signifying improved confidence alignment. However, we notice that there exists a trade-off, as shown in Figure 2 (b), this alignment gain comes at the cost of reduced response accuracy. Such a phenomenon can be attributed to the distinct objectives used for response quality and confidence alignment in the RL fine-tuning process. Fine-tuning a single reward model could fail to evaluate both response quality and alignment performance, inadvertently leading to a heuristic of poor-quality responses with low confidence. In this study, we aim to align confidence with response quality without degrading the quality itself.
@@ 4.1 Dual-component Reward Strategy
To achieve quality and alignment, we decouple the reward function into two component to assess the two objectives, separately:
@@
This strategy is predicated on the notion that accurate assessment of response quality is a prerequisite for aligning it with confidence. Hence, we decouple the evaluation of response quality. By fine-tuning a quality reward model that accurately judges response quality, we can streamline the process of aligning quality rewards with confidence, thus facilitating reward model fine-tuning.
@@
We develop a quality reward model to evaluate the response quality. To train this model, we utilize Reinforcement Learning from Human Feedback (RLHF) datasets, ensuring that the high-quality response receive a higher score than the low-quality ones. We employ a binary ranking loss Touvron et al. (2023) as $\mathcal{L}_{\text{Q}}$:
@@
For confidence alignment, we propose an order-based criterion preserving a consistent order relationship between verbalized confidence and response quality. Specifically, for any pair of samples, $i$ and $j$, an desired relationship between tuples $(x_{i},y_{i},c_{i})$ and $(x_{j},y_{j},c_{j})$ should preserve the order:
@@
The reward function is defined as the sum of the products of pairwise differences in confidence and corresponding reward scores for all samples. This design incentivally rewards the alignment of confidence with the quality of responses, thereby enforcing a direct proportionality between a sample’s stated confidence and its actual quality. It penalizes any deviations from this alignment.
@@
The overall reward function consists of both quality reward and order-preserving alignment reward and can be summarized as follows:
@@ 4.2 RL Fine-tuning LLM
$R(\hat{y}\mid x)$ $=R_{\text{O}}(\hat{y}\mid x)-\beta D_{KL}(\pi_{\theta}(\hat{y}\mid x)\parallel\pi_{0}(\hat{y}\mid x)),$
@@ 4.3 Comparison with PreApproach
We analyze the difference beweteen the PreApproach in Section 3 and CONQORD in Section 4.
@@ 5.1.1 Datasets
Our experiments use two widely-used question-answering (QA) datasets assessing confidence calibration. TruthfulQA Lin et al. (2022b) contains 817 questions spanning 38 categories designed to test language models’ tendency to mimic human falsehoods Zhang et al. (2023a). The Natural Questions (NQ) dataset Kwiatkowski et al. (2019) comprises real anonymized, aggregated queries issued to the Google search engine, forming a question answering dataset. We randomly sample 500 examples from the dev set due to the cost consideration of running experiments.
Dataset Methods ECE $\downarrow$ Pearson Correlation Spearman Correlation Accuracy $\uparrow$
@@
TruthfulQA Vanilla 0.6327 0.0154 0.6611 0.0159 0.6498 0.2387
Top-k 0.5339 -0.0524 0.1347 -0.0577 0.0994 0.3611
CoT+Agg 0.4086 -0.0275 0.5402 -0.0275 0.5402 0.3488
CONQORD 0.1856 0.1086 0.0019 0.1096 0.0017 0.2387
NQ Vanilla 0.4588 0.0786 0.0790 0.0786 0.0790 0.4340
Top-k 0.4046 -0.0268 0.5494 -0.0268 0.5494 0.4940
@@
Table 1: Alignment performance on base model of LLAMA-2 7B on TruthfulQA and NQ datasets. The symbol $\downarrow$ denotes that lower values are preferable, whereas $\uparrow$ indicates that higher values are more desirable.

### v2

@@ Abstract
Despite the success of large language models (LLMs) in natural language generation, much evidence shows that LLMs may produce incorrect or nonsensical text. This limitation highlights the importance of discerning when to trust LLMs, especially in safety-critical domains. Existing methods often express reliability by confidence level, however, their effectiveness is limited by the lack of objective guidance. To address this, we propose CONfidence-Quality-ORDer-preserving alignment approach (CONQORD), which leverages reinforcement learning guided by a tailored dual-component reward function. This function integrates quality reward and order-preserving alignment reward functions. Specifically, the order-preserving reward incentivizes the model to verbalize greater confidence for responses of higher quality to align the order of confidence and quality. Experiments demonstrate that CONQORD significantly improves the alignment performance between confidence and response accuracy, without causing over-cautious. Furthermore, the aligned confidence provided by CONQORD informs when to trust LLMs, and acts as a determinant for initiating the retrieval process of external knowledge. Aligning confidence with response quality ensures more transparent and reliable responses, providing better trustworthiness.
When to Trust LLMs: Aligning Confidence with Response Quality
[…]
@@ 1 Introduction
Large Language Models (LLMs) have excelled in natural language understanding and generation Brown et al. (2020); Anil et al. (2023). However, mounting evidence indicates that LLMs generate incorrect or nonsensical text, including fabricated citations or incorrect medical information, risking errors in critical applications Ji et al. (2023); Zhang et al. (2023b); Agrawal et al. (2023); Du et al. (2023); Cohen et al. (2023). The urgent question is: When can we trust LLMs? Addressing this concern is essential to prevent the uncritical acceptance of misleading information and to guide decisions on when to rely on LLMs versus when to seek external knowledge.
@@
Figure 1: (Top): Vanilla LLMs may generate bad responses and cannot generate confidence. (Middle): Existing methods include verbalizing confidence levels (highlighted in orange) in the output to indicate the model’s uncertainty, yet they may still provide bad responses with overly high confidence, revealing a misalignment between expressed confidence and actual response quality. (Bottom): CONQORD aligns with the confidence and response quality.
Recently, researchers have explored prompting LLMs to output calibrated confidence alongside text Tian et al. (2023) for determining LLMs’ reliability. The essence of confidence calibration is to ensure that the confidence expressed corresponds with the correctness of the response, which is critical for the model’s transparency and trustworthiness. While classification tasks benefit from ground truth labels to calibrate predicted probabilities against actual correctness Guo et al. (2017), generative tasks confront the challenge of calibration without clear ground truths. Current calibration methods for text generation rely on heuristics that consistency among multiple responses Xiong et al. (2023) or the top-k responses facilitate calibration Tian et al. (2023). However, these assumptions often bear little relevance to the intrinsic quality of the responses. Therefore, these strategies frequently result in a misalignment between the expressed confidence and the actual quality of the response, shown in Figure 1. The critical issue is the absence of a gold standard for confidence that directly reflects response quality, leaving models incapable of guiding confidence levels aligned with response quality.
In this study, we explore a Reinforcement Learning (RL) framework to tackle this challenge, designing reward functions that align confidence levels with response quality. This framework takes advantage of the adaptability afforded by the diverse reward functions in RL to bridge the gap between confidence and response quality Guo et al. (2021); Ziegler et al. (2019). A direct reward approach involves rewarding language models for well-aligned response-confidence pairs while penalizing misaligned ones, where the reward model is fine-tuned on the constructed training data. However, this method can inadvertently encourage language models to take shortcuts. Specifically, such a strategy may inadvertently encourage the generation of lower-quality responses paired with correspondingly reduced confidence (empirically demonstrated in Section 3).
To address this issue, it is imperative to devise a reward function that encourages the generation of accurate and well-aligned responses. We propose CONfidence-Quality-ORDer-preserving alignment approach, called CONQORD (sounds as Concord), designing a dual-component reward strategy focusing on response quality and confidence alignment. The reward model comprises:: (i) A quality reward model that rates the response quality. (ii) An order-preserving alignment reward model encourages an ordinal relationship consistency between confidence and quality rating, while penalizing ordinal discrepancies. This order-preserving reward fosters careful self-calibration and adaptability to various contexts. The order-preserving nature of the reward minimizes the impact of outliers. By integrating the quality reward model with the alignment reward function, we apply the Proximal Policy Optimization (PPO) algorithm to harmonize verbalized confidence with response quality, thus preventing the model from becoming cautious.
@@
We conduct experiments using four foundational models, including LLAMA-2 7B, Zephyr 7B, Mistral 7B, and LLAMA-2 13B, across two datasets: NQ and TruthfulQA. Experimental results demonstrate that our CONQORD substantially improves the alignment performance between confidence levels and the quality of responses without inducing excessive caution. Moreover, we evaluate the practicality of CONQORD’s calibrated confidence in adaptive retrieval task Asai et al. (2023); Ding et al. (2024), where confidence scores are used to guide the activation of external knowledge. Our experiments confirm that CONQORD’s confidence alignment reliably dictates the trustworthiness of LLM outputs. CONQORD contributes to making the model-generated responses not only more transparent but also more reliable, through its refined confidence calibration.
@@ 2 Related Works
Confidence, or uncertainty, refers to the degree of certainty or assurance that accompanies a prediction or decision made by a model Geng et al. (2023). The calibration of confidence is essential for the reliability of machine learning systems, as it ensures that predicted probabilities match the true likelihood of outcomes as closely as possible Guo et al. (2017); Minderer et al. (2021). While in traditional classification tasks, this involves aligning predicted probabilities with actual ground truth labels, the task becomes more challenging for generative models due to the inherently ambiguous nature of the ground truth Gawlikowski et al. (2021); Liu et al. (2023).
@@ 3 Confidence Alignment via RL
We adopt a Reinforcement Learning (RL) framework to tackle the challenge of lacking a ground-truth standard for confidence assessment. Different from Supervised Fine-Tuning (SFT), which depends on labeled data, RL offers a more adaptable solution by allowing any indicator as a reward. We follow previous studies Ramamurthy et al. (2023); Wu et al. (2023), regard text language generation as a Markov Decision Process (MDP) while the remaining elements of MDP are listed in Appendix A. In the rest of this section, we elaborate on the reward strategy.
@@ 3.1 Preliminary Alignment Approach
We introduce a preliminary alignment approach (PreApproach) to align confidence with response quality utilizing a reward model. The reward model is fine-tuned on tailored training data consisting of question-response-confidence tuples, to recognize and incentivize confidence alignment.
@@ 3.2 Results and Discussions
The experimental results are illustrated in Figure 2 (a), which shows that PreApproach attains a lower Expected Calibration Error (ECE) Guo et al. (2017) than LLAMA-2 7B, signifying improved confidence alignment.
However, we observe a trade-off, illustrated in Figure 2 (b), where achieving alignment gain is at the expense of diminished response accuracy. This decline may stem from the RL fine-tuning process not focusing on improving response quality, leading to a shortcut where language models tend to produce lower-quality responses with lower confidence. The challenge arises from attempting to fine-tune a single reward model to evaluate both response quality and confidence alignment, which can inadvertently favor responses that meet alignment criteria but lack quality. Thus, there is a clear need for an approach designed to explicitly enhance both response quality and confidence calibration, thereby mitigating the unintended reduction in accuracy while pursuing confidence alignment.
@@ 4.1 Dual-component Reward Strategy
To achieve quality and alignment, we decouple the reward function into two components to assess the two objectives, separately:
@@
This strategy is based on the notion that an accurate assessment of response quality is a prerequisite for aligning it with confidence. Hence, we decouple the evaluation of response quality. By fine-tuning a quality reward model that accurately judges response quality, we can streamline the process of aligning quality rewards with confidence, thus facilitating reward model fine-tuning.
@@
We develop a quality reward model to evaluate the response quality. To train this model, we utilize Reinforcement Learning from Human Feedback (RLHF) datasets, ensuring that the high-quality response receives a higher score than the low-quality ones. We employ a binary ranking loss Touvron et al. (2023) as $\mathcal{L}_{\text{Q}}$:
@@
For confidence alignment, we first introduce an order-based criterion, which preserves a consistent order relationship between verbalized confidence and response quality. Specifically, for any pair of samples, $i$ and $j$, an desired relationship between tuples $(x_{i},y_{i},c_{i})$ and $(x_{j},y_{j},c_{j})$ should preserve the order:
@@
The reward function is defined as the sum of the products of pairwise differences in confidence and corresponding reward scores for all samples. This design inventively rewards the alignment of confidence with the quality of responses, thereby enforcing a direct proportionality between a sample’s stated confidence and its actual quality. It penalizes any deviations from this alignment.
@@
The overall reward function $R_{\text{O}}$ consists of both quality reward $R_{Q}$ and order-preserving alignment reward $R_{A}$, which can be summarized as follows:
@@ 4.2 RL Fine-tuning LLM
$R(\hat{y}\mid x)$ $=R_{\text{O}}(\hat{y}\mid x)$
$-\beta D_{KL}(\pi_{\theta}(\hat{y}\mid x)\parallel\pi_{0}(\hat{y}\mid x)),$
@@ 4.3 Comparison with PreApproach
We analyze the difference between the PreApproach in Section 3 and CONQORD in Section 4.
@@ 5.1.1 Datasets
We conduct experiments on two tasks, hallucination evaluation and question-answering. For hallucination evaluation, we utilize TruthfulQA Lin et al. (2022b), which contains 817 questions spanning 38 categories designed to test language models’ tendency to mimic human falsehoods Zhang et al. (2023a). For question-answering, we adopt the widely-used Natural Questions (NQ) dataset Kwiatkowski et al. (2019), which comprises real anonymized, aggregated queries issued to the Google search engine, forming a question-answering dataset. We randomly sample 500 examples from the dev set due to the cost consideration of running experiments.
Foundation Models Methods ECE $\downarrow$ Pearson Correlation Spearman Correlation Accuracy $\uparrow$
@@
LLAMA-2 7B Vanilla 0.6327 0.0154 6.6×$\times 10^{-1}$ 0.0159 6.5×$\times 10^{-1}$ 0.2387
Top-k 0.5339 -0.0524 1.3×$\times 10^{-1}$0.0577 9.9×$\times 10^{-2}$ 0.3611
CoT+Agg 0.4086 -0.0275 5.4×$\times 10^{-1}$0.0275 5.4×$\times 10^{-1}$ 0.3488
CONQORD 0.1856 0.1086 1.9×$\times 10^{-3}$ 0.1096 1.7×$\times 10^{-3}$ 0.2387
Zephyr 7B Vanilla 0.2132 0.3494 7.2×$\times 10^{-25}$ 0.3814 1.1×$\times 10^{-29}$ 0.4213
Top-k 0.2469 0.2571 8.4×$\times 10^{-14}$ 0.2455 1.1×$\times 10^{-12}$ 0.4419
CoT+Agg 0.2271 0.3952 6.2×$\times 10^{-32}$ 0.4174 8.8×$\times 10^{-36}$ 0.5006
CONQORD 0.1471 0.3992 1.3×$\times 10^{-32}$ 0.4100 1.9×$\times 10^{-34}$ 0.3696
Mistral 7B Vanilla 0.3379 0.0096 7.8×$\times 10^{-1}$ 0.0333 3.4×$\times 10^{-1}$ 0.3244
Top-k 0.2741 0.1531 1.1×${\times}10^{-5}$ 0.1422 4.5×${\times}10^{-5}$ 0.2558
CoT+Agg 0.6021 0.0465 1.8×$\times 10^{-1}$ 0.0411 2.4×$\times 10^{-1}$ 0.2570
CONQORD 0.0228 0.1545 3.3×$\times 10^{-5}$ 0.1509 3.8×$\times 10^{-5}$ 0.3293
LLAMA-2 13B Vanilla 0.5887 0.0578 9.9×${\times}10^{-2}$ 0.0616 7.9×${\times}10^{-2}$ 0.3048
Top-k 0.4950 -0.0296 4.0×${\times}10^{-1}$ 0.0055 8.8×${\times}10^{-1}$ 0.4002
CoT+Agg 0.3696 0.0683 1.3×${\times}10^{-1}$ 0.0683 1.3×${\times}10^{-1}$ 0.5100
CONQORD 0.4942 0.0998 4.3×${\times}10^{-3}$ 0.1789 2.7 ×${\times}10^{-7}$ 0.3011
Table 1: Alignment performance of methods (Vanilla, Top-k, CoT+Agg, and our CONQORD) across the foundation models (LLAMA-2 7B, Zephyr 7B, Mistral 7B, and LLAMA-2 13B) on TruthfulQA dataset. The symbol $\downarrow$ denotes that lower values are preferable, whereas $\uparrow$ indicates that higher values are more desirable.
Foundation Models Methods ECE $\downarrow$ Pearson Correlation Spearman Correlation Accuracy $\uparrow$
LLAMA-2 7B Vanilla 0.4588 0.0786 7.9×$\times 10^{-2}$ 0.0786 7.9×$\times 10^{-2}$ 0.4340
Top-k 0.4046 -0.0268 5.5×$\times 10^{-1}$0.0268 5.5×$\times 10^{-1}$ 0.4940
@@
Zephyr 7B Vanilla 0.3588 0.1481 9.0×$\times 10^{-4}$ 0.1492 8.2×$\times 10^{-4}$ 0.4580
Top-k 0.2746 0.2753 3.8×$\times 10^{-10}$ 0.2833 1.1×${\times}10^{-10}$ 0.3800
CoT+Agg 0.3650 0.1770 6.9×$\times 10^{-5}$ 0.1572 4.2×${\times}10^{-4}$ 0.4360
CONQORD 0.2370 0.2945 4.3×$\times 10^{-10}$ 0.2989 4.6×$\times 10^{-11}$ 0.4500
Mistral 7B Vanilla 0.2258 0.2207 6.3×${\times}10^{-7}$ 0.2197 ${\times}10^{-7}$ 0.3480
Top-k 0.4686 0.1474 1.4×${\times}10^{-3}$ 0.1474 1.4×${\times}10^{-3}$ 0.3780
CoT+Agg 0.3326 0.0576 2.0×${\times}10^{-1}$ 0.0576 2.0×${\times}10^{-1}$ 0.4020
CONQORD 0.0276 0.2435 1.3×${\times}10^{-7}$ 0.2435 1.3×${\times}10^{-7}$ 0.3495
LLAMA-2 13B Vanilla 0.3892 0.0376 4.0×${\times}10^{-1}$ 0.0376 4.0×${\times}10^{-1}$ 0.5040
Top-k 0.3676 0.0898 4.5×${\times}10^{-2}$ 0.0898 4.5×${\times}10^{-2}$ 0.5100
CoT+Agg 0.3110 0.0778 8.2×${\times}10^{-2}$ 0.0778 8.2×${\times}10^{-2}$ 0.5820
CONQORD 0.2922 0.1005 2.5×${\times}10^{-2}$ 0.1160 1.3×${\times}10^{-2}$ 0.4980
Table 2: Alignment performance of methods (Vanilla, Top-k, CoT+Agg, and our CONQORD) across the foundation models (LLAMA-2 7B, Zephyr 7B, Mistral 7B, and LLAMA-2 13B) on NQ dataset. The symbol $\downarrow$ denotes that lower values are preferable, whereas $\uparrow$ indicates that higher values are more desirable.